# Tweets to Citations: Unveiling the Impact of Social Media Influencers on AI Research Visibility

### v1

@@ Appendix A Target-Control Matching
Score: 0.60 |
Target Title: Do Language Models Have Beliefs? Methods for Detecting, Updating, and Visualizing Model Beliefs |
Target Abstract: Do language models have beliefs about the world? Dennett (1995) famously argues that even thermostats have beliefs, on the view that a bel… |
Control Title: An Interpretability Illusion for BERT |
Control Abstract: We describe an"interpretability illusion"that arises when analyzing the BERT model. Activations of individual neurons in the network… |
Score: 0.68 |
Target Title: I2D2: Inductive Knowledge Distillation with NeuroLogic and Self-Imitation |
Target Abstract: Commonsense capabilities of pre-trained language models dramatically improve with scale, leading many to believe that scale is the… |
Control Title: Do language models have coherent mental models of everyday things? |
Control Abstract: When people think of everyday things like an egg, they typically have a mental image associated with it. This allows them to correctly judge… |
Score: 0.77 |
Target Title: MViTv2: Improved Multiscale Vision Transformers for Classification and Detection |
Target Abstract: In this paper, we study Multiscale Vision Transformers (MViTv2) as a unified architecture for image and video classification, as… |
Control Title: Object-Region Video Transformers |
Control Abstract: Recently, video transformers have shown great success in video understanding, exceeding CNN performance; yet existing video trans… |
Score: 0.78 |
Target Title: 3D Clothed Human Reconstruction in the Wild |
Target Abstract: . Although much progress has been made in 3D clothed human reconstruction, most of the existing methods fail to produce robust results… |
Control Title: UNIF: United Neural Implicit Functions for Clothed Human Reconstruction and Animation |
Control Abstract: We propose united implicit functions (UNIF), a part-based method for clothed human reconstruction and animation with raw scans and… |
Score: 0.80 |
Target Title: Open-Source Large Language Models Outperform Crowd Workers and Approach ChatGPT in Text-Annotation Tasks |
Target Abstract: This study examines the performance of open-source Large Language Models (LLMs) in text annotation tasks and compares it with propr… |
Control Title: Can ChatGPT Reproduce Human-Generated Labels? A Study of Social Computing Tasks |
Control Abstract: The release of ChatGPT has uncovered a range of possibilities whereby large language models (LLMs) can substitute human intelligence… |
Score: 0.85 |
Target Title: Leveraging Knowledge in Multilingual Commonsense Reasoning |
Target Abstract: Commonsense reasoning (CSR) requires models to be equipped with general world knowledge. While CSR is a language-agnostic process… |
Control Title: It’s All in the Heads: Using Attention Heads as a Baseline for Cross-Lingual Transfer in Commonsense Reasoning |
Control Abstract: Commonsense reasoning is one of the key problems in natural language processing, but the relative scarcity of labeled data holds back the prog… |
@@ Appendix B Additional Analysis Plots
Appendix C Limitations
@@
Binary Gender Analysis: Our study utilized a gender prediction API that only outputs binary genders (male and female), limiting our ability to capture the full spectrum of gender identities. This binary approach oversimplifies the complex nature of gender and excludes non-binary, transgender, and gender-nonconforming individuals. Consequently, our findings may not accurately represent the diversity of gender identities within the ML community.
Accuracy with Non-Western Names: The gender prediction API may exhibit lower accuracy when analyzing non-Western names. This limitation potentially introduces bias in our gender distribution analysis, particularly in underrepresenting or misrepresenting authors from diverse cultural and ethnic backgrounds. As a result, the study’s conclusions regarding gender disparity might not fully capture the global diversity of the ML community.
Self-Reported Affiliations: The affiliations in our dataset are self-reported, which can lead to inconsistencies such as misspellings or missing location information. This limitation affects the accuracy of our geographical analysis, as the data may not accurately reflect the true affiliations or locations of the authors. Consequently, our observations regarding geographical concentration and diversity must be interpreted with caution.
Lack of Randomized Control Trial: The study did not include a randomized control trial, often considered the gold standard for establishing causality. Without this, we cannot conclusively determine whether the patterns observed are directly attributable to the influencers’ activities or are coincidental. This limitation means that our findings, while suggestive of trends, cannot definitively establish cause-and-effect relationships between influencer endorsements and paper visibility or citation counts.
In summary, these limitations highlight the need for cautious interpretation of our findings. Future research should consider incorporating more inclusive gender identification methods, improving the accuracy of cultural and geographical representation, and employing more rigorous experimental designs, such as randomized control trials, to strengthen the conclusions drawn from such studies.

### v2

@@ Appendix A Target-Control Matching
Qualitatively, the matched pairs are very similar in topic, almost always covering the same sub-field of research (for example, language model hallucinations). We supply a random sample of matched pairs in table 6. The distribution of scores is shown in table 9, and shows that the matched pairs have high cosine similarity scores. Our results are consistent across many choices of matching schemes, such as when we use more than three quantiles for binned variables, or exclude author characteristics altogether.
Score: 0.09 |
Target Title: Unifying Vision, Text, and Layout for Universal Document Processing |
Target Abstract: We propose Universal Document Processing (UDOP), a foundation Document AI model which unifies text, image, and layout modalities together with varied task forma… |
Control Title: SwinTextSpotter: Scene Text Spotting via Better Synergy between Text Detection and Text Recognition |
Control Abstract: End-to-end scene text spotting has attracted great attention in recent years due to the success of excavating the intrinsic synergy of the scene text detection … |
Score: 0.09 |
Target Title: reStructured Pre-training |
Target Abstract: In this work, we try to decipher the internal connection of NLP technology development in the past decades, searching for essence, which rewards us with a (pote… |
Control Title: Adapting BigScience Multilingual Model to Unseen Languages |
Control Abstract: We benchmark different strategies of adding new languages (German and Korean) into the BigScience’s pretrained multilingual language model with 1.3 billion para… |
Score: 0.12 |
Target Title: Quantifying Memorization Across Neural Language Models |
Target Abstract: Large language models (LMs) have been shown to memorize parts of their training data, and when prompted appropriately, they will emit the memorized training dat… |
Control Title: How Much Data Are Augmentations Worth? An Investigation into Scaling Laws, Invariance, and Implicit Regularization |
Control Abstract: Despite the clear performance benefits of data augmentations, little is known about why they are so effective. In this paper, we disentangle several key mechani… |
Score: 0.07 |
Target Title: Do Language Models Know When They’re Hallucinating References? |
Target Abstract: State-of-the-art language models (LMs) are famous for"hallucinating"references. These fabricated article and book titles lead to harms, obstacles to their use, … |
Control Title: Retrieving Supporting Evidence for LLMs Generated Answers |
Control Abstract: Current large language models (LLMs) can exhibit near-human levels of performance on many natural language tasks, including open-domain question answering. Unfo… |
Score: 0.10 |
Target Title: Singularity: Planet-Scale, Preemptive and Elastic Scheduling of AI Workloads |
Target Abstract: Lowering costs by driving high utilization across deep learning workloads is a crucial lever for cloud providers. We present Singularity, Microsoft’s globally d… |
Control Title: Edge Impulse: An MLOps Platform for Tiny Machine Learning |
Control Abstract: Edge Impulse is a cloud-based machine learning operations (MLOps) platform for developing embedded and edge ML (TinyML) systems that can be deployed to a wide r… |
Score: 0.09 |
Target Title: A Large-Scale Study on Unsupervised Spatiotemporal Representation Learning |
Target Abstract: We present a large-scale study on unsupervised spatiotemporal representation learning from videos. With a unified perspective on four recent image-based framewo… |
Control Title: Target-Aware Object Discovery and Association for Unsupervised Video Multi-Object Segmentation |
Control Abstract: This paper addresses the task of unsupervised video multi-object segmentation. Current approaches follow a two-stage paradigm: 1) detect object proposals using … |
@@ Appendix B Additional Analysis Plots
Appendix C Causal Inference
For our causal inference analysis, we draw primarily from the model, techniques, and assumptions presented in Elazar et al. [2023]. In turn, they point interested readers to Imbens and Rubin [2015] and Feder et al. [2021].
Background.
In our work, we study the outcome as a binary variable , indicating if a paper is "highly-cited" () or "less-cited" (). By introducing our treatment , we can define two cases for any given paper: , the outcome if the paper were shared by influencers (when ), or , the outcome if the paper were not shared by influencers (when ). In actuality, we can only observe exactly one of these for each paper. Rather, the goal of causal inference is to estimate the change in outcome if the treatment were flipped (Figure 12). Specifially, we hope to answer the question, does influencer sharing increase the likelihood of a paper being highly-cited?
To accomplish this, we use the average treatment effect on treated:
which is the expected increase in highly-cited papers when shared by influencers, for those already shared.
Negative Outcome Control and Difference-in-Difference.
To debias the effects of our unobserved confounders, we will utilize the Negative Outcome Control (NOC) framework Card and Krueger [1993], Lipsitch et al. [2010]. Specifically, we will use Difference-in-Difference (DiD) to estimate ATET over two time periods Sheppard et al. [1999]:
which can be estimated by a fitting a logistic model and bootstrapping for confidence intervals Elazar et al. [2023].
Sofer et al. [2016] establishes that NCO and DiD-based estimators are equivalent, meaning ATET can be rewritten as:
and solved by taking the difference across treatment groups of the difference between and .
Assumptions.
To perform our causal estimate, we assume the following conditions hold:
1.
Ignorability:
2.
Positivity:
3.
Consistency:
4.
Negative control:
5.
Additive Equi-confounding:
In other words, ignorability assumes are the only confounders with an effect on the outcome; positivity means each paper has a non-zero chance of being and not being influencer-shared; consistency establishes that the potential outcomes and observed outcomes agree for a given treatment level ; negative control assumes that the NCO is not causally affected by the treatment ; additive equi-confounding relies on a linear (or logistic) outcome model for .
Review Scores as a Negative Control Outcome.
The goal of the Negative Outcome Control (NOC) framework Card and Krueger [1993], Lipsitch et al. [2010] is to find a Negative Control Outcome (NCO) variable that is affected by the same confounders as the outcome , but not the treatment. This way, we can de-bias the effects of confounders not captured by the metadata and text embeddings collected from S2. For this, we choose the average conference review score for each paper, collected as described in Section 3. We believe that this sub-sample is representative for the following reasons: (1) influencers often share preprint works before conference acceptance or proceedings are released; (2) our matching accounts for publication venue as a confounder; (3) using bootstrapping on the three tests from Table 5, we find there is not significant evidence to reject that the sub-sample and full dataset citations come from different distributions (-value ).
Matching and Causal Inference.
Elazar et al. [2023] use tripartite matching on a number of covariates Zhang et al. [2021], and only sampled from one venue. To control the topic covariate, they use SPECTER embeddings to build 20 topic clusters and minimize distance for numerical variables (author citations, institution rank, etc.)
Instead, we group the treated and control samples based on a mix of exact matching (venue, year, open access) and quantile binning (number of authors, h-index, and author cites). From these groups, we match samples by minimizing the difference of text embeddings. This way, we can most accurately capture increased nuance between "viral" topics, techniques, and models.
Additionally, we use first-author information, rather than both minimum and average, to increase our group sizes and improve the quality of topic matches while maintaining descriptive author details.
Appendix D Limitations
@@
Binary Gender Analysis: Our study utilized a gender prediction API that only outputs binary genders (male and female), limiting our ability to capture the full spectrum of gender identities. Consequently, our findings may not accurately represent the diversity of gender identities within the ML community.
Accuracy with Non-Western Names: The gender prediction API may exhibit lower accuracy when analyzing non-Western names. As a result, the study’s conclusions regarding gender disparity might not fully capture the global diversity of our samples.
Self-Reported Affiliations: The affiliations in our dataset are self-reported, which can lead to inconsistencies such as misspellings or missing location information. This limitation affects the accuracy of our geographical analysis, as the data may not accurately reflect the true affiliations or locations of the authors. Consequently, our observations regarding geographical concentration and diversity must be interpreted with caution. Additionally, this limitation posed as a barrier to investigating institutions as an observed confounder for causal inference. We invite future studies with more complete affiliation data to investigate this.
Lack of Randomized Control Trial: The study did not include a randomized control trial, often considered the gold standard for establishing causality. Without this, we cannot conclusively determine whether the patterns observed are directly attributable to the influencers’ activities or are coincidental. Though we attempt to remedy this through causal inference, such techniques are contingent upon the validity of our assumptions.